# PIS3R: Very Large Parallax Image Stitching via Deep 3D Reconstruction

### v1

@@ Abstract
Image stitching aim to align two images taken from different viewpoints into one seamless, wider image. However, when the 3D scene contains depth variations and the camera baseline is significant, noticeable parallax occurs—meaning the relative positions of scene elements differ substantially between views. Most existing stitching methods struggle to handle such images with large parallax effectively. To address this challenge, in this paper, we propose an image stitching solution called PIS3R that is robust to very large parallax based on the novel concept of deep 3D reconstruction. First, we apply visual geometry grounded transformer to two input images with very large parallax to obtain both intrinsic and extrinsic parameters, as well as the dense 3D scene reconstruction. Subsequently, we reproject reconstructed dense point cloud onto a designated reference view using the recovered camera parameters, achieving pixel-wise alignment and generating an initial stitched image. Finally, to further address potential artifacts such as holes or noise in the initial stitching, we propose a point-conditioned image diffusion module to obtain the refined result. Compared with existing methods, our solution is very large parallax tolerant and also provides results that fully preserve the geometric integrity of all pixels in the 3D photogrammetric context, enabling direct applicability to downstream 3D vision tasks such as SfM. Experimental results demonstrate that the proposed algorithm provides accurate stitching results for images with very large parallax, and outperforms the existing methods qualitatively and quantitatively.
Figure 1: PIS3R for very large parallax image stitching. (a) The image stitching task can be classified (Sec. B in Appendix) into three cases based on the baseline between target images: pure rotation, slight parallax, and very large parallax. In this paper, we propose a novel image stitching solution named PIS3R, specifically for stitching images with very large parallax. (b) Representative existing image stitching methods (Zaragoza et al. 2014; Li et al. 2018; Nie et al. 2023) are unable to handle stitching tasks involving very large parallax. Instead, our solution is achieving promising results in such challenging scenario.
* Equal contribution. Corresponding author: yizhenlao@hnu.edu.cnProject page: https://github.com/zmhcasmy/PIS3R
1 Introduction
Image stitching is a useful technique used to generate a wide field-of-view (FoV) scene by combining multiple images captured with limited FoV. It plays a vital role across various domains, including autonomous driving, medical imaging, surveillance, and virtual reality.
@@ 1 Introduction
Image warping is a key challenge in image stitching. Homography, a widely used planar transformation model, assumes a flat scene (Hartley and Zisserman 2003). However, in real-world scenes with depth variation and large camera baselines, this assumption fails, leading to parallax effects and noticeable misalignments—especially near object boundaries.
1.1 Motivations
Very large parallax. To mitigate parallax artifacts in image stitching, adaptive warping methods have been developed that divide images into grids or pixels and apply locally varying transformations (Gao, Kim, and Brown 2011; Lee and Sim 2018; Li et al. 2018; Lin et al. 2011; Zaragoza et al. 2014; Zhang et al. 2016). Energy minimization frameworks further optimize these warps to reduce geometric distortion (Li et al. 2018; Lin et al. 2011; Zhang et al. 2016). Some approaches employ seam-cutting techniques to align local regions and conceal misalignments in others (Gao et al. 2013; Lin et al. 2016; Zhang and Liu 2014). However, existing smooth warping methods (Gao, Kim, and Brown 2011; Li et al. 2018; Lin et al. 2011; Zaragoza et al. 2014; Zhang et al. 2016) often fail under very large parallax, where adjacent pixels in one image may not correspond to nearby pixels in the other view. While video-based methods using epipolar geometry have been proposed (Lee and Sim 2018), they rely on temporal motion cues and are not directly applicable to still image stitching.
@@ 1.1 Motivations
In summary, if we can stitch images with very large parallax and produce not only visually seamless result but also are capable to support downstream 3DV tasks.
1.2 Contribution
As shown in Fig. 2, to tackle the challenges of existing methods in handling extreme parallax, we introduce a novel image stitching framework named PIS3R, which leverages deep 3D reconstruction to achieve robust alignment under very large parallax conditions. Our method begins by applying a feed-forward scene reconstruction approach VGGT (Wang et al. 2025) to a pair of uncalibrated input images, recovering both intrinsic and extrinsic camera parameters along with a dense 3D reconstruction of the scene. The resulting point cloud is then reprojected onto a designated reference view using the estimated camera poses, yielding an initial stitched image with pixel-wise geometric alignment. To further refine this result and mitigate artifacts such as holes or noise introduced during reprojection, we incorporate a point-conditioned image diffusion module.
Unlike existing approaches that primarily focus on visual seamlessness, our method explicitly preserves the 3D geometric consistency of every pixel in the photogrammetric space. This not only enables high-quality stitching under large parallax but also ensures compatibility with downstream 3D vision tasks such as SfM, SLAM, and depth estimation. Experimental results confirm that PIS3R outperforms both qualitatively and quantitatively compared to existing methods. Our contributions center around:
  • We propose a novel pipeline for stitching images with very large parallax based on the deep 3D reconstruction results and followed by a diffusion-based refinement, called PIS3R, which is a fundamental departure from the technical pipeline of existing methods.
@@ 2 Related Work
Traditional image stitching. 1) Adaptive warping. Traditional stitching methods often leverage discriminative features like SIFT (Lowe 2004) for estimating flexible warps such as DHW (Gao, Kim, and Brown 2011), APAP (Zaragoza et al. 2014), and TFA (Lin et al. 2011). Extensions like DFW (Li et al. 2017) and GES-GSP (Lin et al. 2021a) incorporate lines, edges, or semantic planes to enhance structural preservation. Furthermore, depth and planar priors have been utilized to improve alignment under challenging conditions (Liu et al. 2020; Zhou et al. 2018). 2) Seam cutting. Seam-cutting is commonly employed as a post-processing step for artifact reduction. Methods define energy terms based on photometric, gradient, or motion cues (Kwatra et al. 2003; Agarwala et al. 2004; Bao, Sun, and Yang 2013), and minimize them via graph cuts to produce visually coherent results. Seam-based alignment optimization has also been explored for minimizing stitching distortions (Zhang and Liu 2014; Lin et al. 2016; Zhou, Zhang, and Bao 2018). However, these geometry-driven methods face two limitations: 1) they are sensitive to missing or weak structural features, leading to alignment failure; and 2) when geometric features are abundant, the computational overhead increases significantly.
Deep image stitching. These approaches learn high-level semantic representations from data. They have been developed under supervised (Nie et al. 2023; Lin et al. 2021b; Zhou, Zhang, and Bao 2022; Huang, Wang, and Fu 2021; Zhou et al. 2022), weakly supervised (Nie, Zhang, and Fang 2022), and unsupervised paradigms (Nie, Zhang, and Fang 2021), demonstrating strong adaptability to complex scenes. Among these, unsupervised methods are particularly appealing due to the scarcity of ground truth stitched images. However, most of them rely on homography-based alignment, which struggles under very large parallax as shown in Fig. 2, often resulting in visible blur and artifacts in the reconstructed regions without mentioning being unable to preserve the 3D projection consistency.
3 Methodology
@@ 3.1 PIS3R Pipeline
Step 1: 3D reconstruction. Given a sparse set of two or more images, we first establish their point cloud representation using a dense stereo model, as illustrated in Sec. 3.2.
Step 2: Point Cloud Reprojection. Subsequently, we reproject the point cloud onto the input camera poses predicted by the dense stereo model to obtain a preliminary stitched image, as illustrated in Sec. 3.3.
Step 3: Refinement of Reprojected Images. To achieve visually pleasing results while preserving the geometric structure of the scene , we train a point-conditioned image diffusion model to refine the noisy preliminary stitched image, as depicted in Sec. 3.4.
3.2 Deep-based 3D Reconstruction
As shown in Fig. 3, we first construct a point cloud representation from the input images to obtain information about the entire scene. Certainly different from prior approaches that achieve image stitching through geometry-based 3D reconstruction, we adopts a deep-based reconstruction framework. Specifically, we employ VGGT (Wang et al. 2025) [^1] https://github.com/facebookresearch/vggtto reconstruct the entire scene and estimate camera parameters simultaneously. The input consists of a sequence of RGB images capturing the same 3D scene $\left(I_{i}\right)_{i=1}^{N}{\in\mathbb{R}^{{H}\times W\times 3}}$ to corresponding 3D annotation information via a transformer, which can be described as:
@@ 3.2 Deep-based 3D Reconstruction
As shown in Fig. 4, we compare the point cloud representations generated by the geometry-based reconstruction framework COLMAP  (Schönberger et al. 2016; Schönberger and Frahm 2016) [^2] https://github.com/colmap/colmap and VGGT under three different input settings: pure rotation, slight parallax, and very large parallax. The reconstruction performance of COLMAP under sparse viewpoints is predictable, while the reconstruction results of VGGT are remarkable. The quality of the point cloud and the accuracy of the predicted camera poses are crucial for image stitching using 3D reconstruction. This directly determines the geometric structure and quality of final stitched image.
3.3 Image Stitching via Point Cloud Reprojection
As shown in Fig. 3, to obtain the reprojected image $\hat{S}_{i}$ from $I_{i}$, we first generate the transformed point cloud $P_{c}^{i}\in\mathbb{R}^{NHW\times 3}$ by converting the predicted point cloud $P_{w}\in\mathbb{R}^{NHW\times 3}$ from the world coordinate to the camera coordinate $O_{c}^{i}$ using the camera extrinsic parameters (rotation $R^{i}$ and translation $t^{i}$). The transformation from world coordinates to camera coordinates can be described as:
@@ 3.3 Image Stitching via Point Cloud Reprojection
$P_{uv}^{i}=\frac{K_{i}P_{c}^{i}}{Z},$ (3)
where $K_{i}$ is the intrinsic matrix of camera $i$, and Z{Z} denotes the depth of each 3D point in $P^{i}_{c}$. Finally, the reprojected image $\hat{S}_{i}$ is constructed by normalizing the pixel values at the reprojected positions according to Eq. 4:
$\hat{S}_{i}=Normalize\left(P_{uv}^{i}\right),$ (4)
where $Normalize(\cdot)$ denotes normalize the reprojected pixel values to the non-negative coordinate space.
We integrate the point map with its corresponding RGB image to obtain the reprojected image at the respective pose. This provides coarse 3D scene information for image stitching. However, sparse inputs offer insufficient 3D cues, which cause significant missing regions, occlusions, and artifacts in the reconstructed point cloud. These issues further cause missing regions and pixel misalignments in the reprojected images. Therefore, we propose integrating image diffusion models to obtain visually pleasing stitched image.
3.4 Diffusion-based Stitching Refinement
@@
Figure 5: Qualitative comparison of different image restoration diffusion models. We evaluated the performance of several state-of-the-art image restoration diffusion models in refining reprojected images, including DiffIR  (Xia et al. 2023), DDRM (Kawar et al. 2022), and RDDM  (Liu et al. 2024).
3.5 Implementation

### v2

@@ Abstract
Image stitching aim to align two images taken from different viewpoints into one seamless, wider image. However, when the 3D scene contains depth variations and the camera baseline is significant, noticeable parallax occurs—meaning the relative positions of scene elements differ substantially between views. Most existing stitching methods struggle to handle such images with large parallax effectively. To address this challenge, in this paper, we propose an image stitching solution called PIS3R that is robust to very large parallax(VLP) based on the novel concept of deep 3D reconstruction. First, we apply visual geometry grounded transformer to two input images with VLP to obtain both intrinsic and extrinsic parameters, as well as the dense 3D scene reconstruction. Subsequently, we reproject reconstructed dense point cloud onto a designated reference view using the recovered camera parameters, achieving pixel-wise alignment and generating an initial stitched image. Finally, to further address potential artifacts such as holes or noise in the initial stitching, we propose a point-conditioned image diffusion module to obtain the refined result. Compared with existing methods, our solution is VLP tolerant and also provides results that fully preserve the geometric integrity of all pixels in the 3D photogrammetric context, enabling direct applicability to downstream 3D vision tasks such as SfM. Experimental results demonstrate that the proposed algorithm provides accurate stitching results for images with VLP, and outperforms the existing methods qualitatively and quantitatively.
Figure 1: PIS3R for very large parallax image stitching. (a) The image stitching task can be classified (Sec. B in Appendix) into three cases based on the baseline between target images: pure rotation, slight parallax, and very large parallax. In this paper, we propose a novel image stitching solution named PIS3R, specifically for stitching images with very large parallax. (b) Representative existing image stitching methods [apap_zaragoza2014projective, ELA_li2018parallax, udis2_nie2023parallax] are unable to handle stitching tasks involving very large parallax. Instead, our solution is achieving promising results in such challenging scenario.
I Introduction
TABLE I: Comparison of the proposed method PIS3R, vs. the state-of-the-art parallax-tolerate solutions.
Method Adaptive Warping Seam Driven Deep Learning
APAP[apap_zaragoza2014projective] DFW[DFW_li2017dual] ELA[ELA_li2018parallax] OBJ-GSP[obj-gsp_cai2025object] PTIS[PTIS_Zhang_2014_CVPR] SEAGULL[lin2016seagull] UDIS[udis_d_nie2021unsupervised] UDIS++[udis2_nie2023parallax] PIS3R
TPAMI’14 ICIP’17 TMM’18 AAAI’25 CVPR’14 ECCV’16 TIP’21 ICCV’23
Pure rotation
Slight parallax
Very large parallax
Image stitching is a useful technique used to generate a wide field-of-view (FoV) scene by combining multiple images captured with limited FoV. It plays a vital role across various domains, including autonomous driving [auto_driveing_kinzig2022real], medical imaging, surveillance, and virtual reality [VR_anderson2016jump].
@@ I Introduction
Image warping is a key challenge in image stitching. Homography, a widely used planar transformation model, assumes a flat scene [hartley2003multiple]. However, in real-world scenes with depth variation and large camera baselines, this assumption fails, leading to parallax effects and noticeable misalignments—especially near object boundaries.
I-A Motivations
Very large parallax (VLP). To mitigate parallax artifacts in image stitching, adaptive warping methods have been developed that divide images into grids or pixels and apply locally varying transformations [gao2011dual, lee2018video, ELA_li2018parallax, lin2011smoothly, apap_zaragoza2014projective, zhang2016multi]. Energy minimization frameworks further optimize these warps to reduce geometric distortion [ELA_li2018parallax, lin2011smoothly, zhang2016multi]. Some approaches employ seam-cutting techniques to align local regions and conceal misalignments in others [gao2013seam, lin2016seagull, zhang2014parallax]. However, existing smooth warping methods [gao2011dual, ELA_li2018parallax, lin2011smoothly, apap_zaragoza2014projective, zhang2016multi] often fail under VLP, where adjacent pixels in one image may not correspond to nearby pixels in the other view. While video-based methods using epipolar geometry have been proposed [lee2018video], they rely on temporal motion cues and are not directly applicable to still image stitching.
@@ I-A Motivations
In summary, if we can stitch images with VLP and produce not only visually seamless result but also are capable to support downstream 3DV tasks.
I-B Contribution
As shown in Fig. 2 and Tab. I, to address the limitations of existing methods under extreme parallax, we propose PIS3R, a novel image stitching framework leveraging deep 3D reconstruction for robust alignment in very large parallax (VLP) scenarios. Our method first applies the feed-forward scene reconstruction model VGGT [wang2025vggt] to uncalibrated image pairs, jointly recovering camera intrinsics, extrinsics, and a dense 3D scene. The reconstructed point cloud is then reprojected onto a reference view using the estimated poses, producing an initial stitched image with pixel-wise geometric alignment. A point-conditioned diffusion module further refines this result, filling holes and suppressing reprojection artifacts.
Unlike prior work that prioritizes visual seamlessness, PIS3R explicitly preserves 3D photogrammetric consistency for every pixel, enabling high-quality stitching under large parallax and ensuring compatibility with downstream 3D vision tasks like SfM, SLAM, and depth estimation. Experiments show PIS3R outperforms existing methods both qualitatively and quantitatively. Our contributions are:
  • We propose a novel pipeline for stitching images with VLP based on the deep 3D reconstruction results and followed by a diffusion-based refinement, called PIS3R, which is a fundamental departure from the technical pipeline of existing methods.
@@ II Related Work
Traditional image stitching. 1) Adaptive warping. Traditional stitching methods often leverage discriminative features like SIFT [lowe2004distinctive] for estimating flexible warps such as DHW [gao2011dual], APAP [apap_zaragoza2014projective], and TFA [lin2011smoothly]. Extensions like DFW [DFW_li2017dual], GES-GSP [lin2021gesgsp] and OBJ-GSP [obj-gsp_cai2025object] incorporate lines, edges, mesh or semantic planes to enhance structural preservation. Furthermore, depth and planar priors have been utilized to improve alignment under challenging conditions [liu2020attentional, zhou2018stereo]. 2) Seam cutting. Seam-cutting is commonly employed as a post-processing step for artifact reduction. Methods define energy terms based on photometric, gradient, or motion cues [kwatra2003graphcut, agarwala2004interactive, bao2013fast], and minimize them via graph cuts to produce visually coherent results. Seam-based alignment optimization has also been explored for minimizing stitching distortions [zhang2014parallax, lin2016seagull, zhou2018stitchability]. However, these geometry-driven methods face two limitations: 1) they are sensitive to missing or weak structural features, leading to alignment failure; and 2) when geometric features are abundant, the computational overhead increases significantly.
Deep image stitching. These approaches learn high-level semantic representations from data. They have been developed under supervised [udis2_nie2023parallax, lin2021deep, zhou2022learning, huang2021image, zhou2022warpstitch], weakly supervised [nie2022unsupervised], and unsupervised paradigms [nie2021unsupervised], demonstrating strong adaptability to complex scenes. Among these, unsupervised methods are particularly appealing due to the scarcity of ground truth stitched images. However, most of them rely on homography-based alignment, which struggles under very large parallax as shown in Fig. 2, often resulting in visible blur and artifacts in the reconstructed regions without mentioning being unable to preserve the 3D projection consistency.
III Methodology
@@ III-A PIS3R Pipeline
Step 1: 3D reconstruction. Given a sparse set of two or more images, we first establish their point cloud representation using a dense stereo model, as illustrated in Sec. III-B.
Step 2: Point Cloud Reprojection. Subsequently, we reproject the point cloud onto the input camera poses predicted by the dense stereo model to obtain a preliminary stitched image, as illustrated in Sec. III-C.
Step 3: Refinement of Reprojected Images. To achieve visually pleasing results while preserving the geometric structure of the scene , we train a point-conditioned image diffusion model to refine the noisy preliminary stitched image, as depicted in Sec. III-D.
III-B Deep-based 3D Reconstruction
As shown in Fig. 3, we first construct a point cloud representation from the input images to obtain information about the entire scene. Certainly different from prior approaches that achieve image stitching through geometry-based 3D reconstruction, we adopts a deep-based reconstruction framework. Specifically, we employ VGGT [wang2025vggt] [^1] https://github.com/facebookresearch/vggtto reconstruct the entire scene and estimate camera parameters simultaneously. The input consists of a sequence of RGB images capturing the same 3D scene $\left(I_{i}\right)_{i=1}^{N}{\in\mathbb{R}^{{H}\times W\times 3}}$ to corresponding 3D annotation information via a transformer, which can be described as:
@@ III-B Deep-based 3D Reconstruction
As shown in Fig. 4, we compare the point cloud representations generated by the geometry-based reconstruction framework COLMAP  [colmap_schoenberger2016mvs, colmap_schoenberger2016sfm] [^2] https://github.com/colmap/colmap and VGGT under three different input settings: pure rotation, slight parallax, and very large parallax. The reconstruction performance of COLMAP under sparse viewpoints is predictable, while the reconstruction results of VGGT are remarkable. The quality of the point cloud and the accuracy of the predicted camera poses are crucial for image stitching using 3D reconstruction. This directly determines the geometric structure and quality of final stitched image.
III-C Image Stitching via Point Cloud Reprojection
As shown in Fig. 3, to obtain the reprojected image $\hat{S}_{i}$ from all input images, we first generate the transformed point cloud $P_{c}^{i}\in\mathbb{R}^{NHW\times 3}$ by converting the predicted point cloud $P_{w}\in\mathbb{R}^{NHW\times 3}$ from the world coordinate to the camera coordinate $O_{c}^{i}$ using the camera extrinsic parameters (rotation $R^{i}$ and translation $t^{i}$). The transformation from world coordinates to camera coordinates can be described as:
@@ III-C Image Stitching via Point Cloud Reprojection
$P_{uv}^{i}=\frac{K_{i}P_{c}^{i}}{Z_{c}^{i}},$ (3)
where $K_{i}$ is the intrinsic matrix of camera $i$, and $Z_{c}^{i}$} denotes the depth of 3D point in $P^{i}_{c}$.
Subsequently, we construct a pixel grid $\mathcal{P}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times 4}$, whose spatial resolution $(H_{i},W_{i})$ is dynamically determined from the coordinate range of the projected points $P_{uv}^{i}$. For each integer pixel location $(x,y)$, we define the set of 3D points that project to it as:
$\mathcal{N}(x,y)=\left\{j\;\middle|\;\operatorname{round}\big(P_{uv}^{i}(j)\big)=(x,y)\right\}.$ (4)
Among these, we select the point with minimum depth in the camera coordinate system:
$k=\arg\min_{j\in\mathcal{N}(x,y)}Z_{c}^{i}(j),$ (5)
$\mathcal{P}_{i}(x,y)=\mathcal{G}\big(c_{k},\,Z_{c}^{i}(k)\big),$
where $P_{w}(j)$ denotes the $j$-th 3D point in world coordinates, $Z_{c}^{i}(j)$ its depth in the $i$-th camera frame, and $P_{uv}^{i}(j)$ its 2D projection onto the corresponding image plane in pixel coordinates.
To alleviate the burden on the diffusion model during refinement, we further process the holes within the projected regions. Let $S_{i}$ denote the set of all filled pixel positions. We then compute the convex hull of $S_{i}$, denoted $\mathcal{H}_{i}$, which represents the minimal convex region that encompasses all reprojected pixels:
$S_{i}=\left\{(x,y)\in\mathbb{Z}^{2}\,\middle|\,\exists\,j:\ \operatorname{round}\!\big(P_{uv}^{i}(j)\big)=(x,y)\right\}.$ (6)
$\mathcal{H}_{i}=\mathrm{Convex\ Hull}(S_{i}).$ (7)
Based on $S_{i}$ and $\mathcal{H}_{i}$, we construct three binary masks on the integer lattice:
$M_{\mathrm{f}}(x,y)$ $=\mathbf{1}\!\big[(x,y)\in S_{i}\big],$ (8)
$M_{\mathrm{r}}(x,y)$ $=\mathbf{1}\!\big[(x,y)\in\mathcal{H}_{i}\big],$
$M_{\mathrm{h}}(x,y)$ $=M_{\mathrm{r}}(x,y)\big(1-M_{\mathrm{f}}(x,y)\big),$
where, $M_{\mathrm{f}}$ indicates the originally observed (filled) pixels, $M_{\mathrm{r}}$ covers all integer pixels lying inside or on the boundary of the convex hull, and $M_{\mathrm{h}}$ identifies interior hole pixels—those within $\mathcal{H}_{i}$ but not in $S_{i}$.
For any hole pixel $(x,y)$ with $M_{\mathrm{h}}(x,y)=1$, we define a neighborhood $\mathcal{N}_{K}(x,y)$ consisting of the $K$ nearest filled pixels within the hull:
$\mathcal{N}_{K}(x,y)\subseteq\left\{(u,v)\,\middle|\,M_{\mathrm{f}}(u,v)=1,\ M_{\mathrm{r}}(u,v)=1\right\}.$ (9)
Each neighbor is assigned a Gaussian weight based on its Euclidean distance to $(x,y)$:
$\alpha\!\big((x,y),(u,v)\big)=\exp\!\left(-\frac{\|(x-u,\ y-v)\|_{2}^{2}}{2\sigma^{2}}\right).$ (10)
The hole color $\mathbf{H}_{i}(x,y)$ is then computed as a distance-weighted average of the RGB values of these neighbors:
$\mathbf{H}_{i}(x,y)=\frac{\sum_{(u,v)\in\mathcal{N}_{K}(x,y)}\alpha\!\big((x,y),(u,v)\big)\,\mathbf{C}_{i}(u,v)}{\sum_{(u,v)\in\mathcal{N}_{K}(x,y)}\alpha\!\big((x,y),(u,v)\big)},$ (11)
where $\mathbf{C}_{i}(u,v)\in\mathbb{R}^{3}$ denotes the RGB color corresponding to the 3D point.
Finally, the complete reprojected image $\hat{S}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times 3}$ is obtained by rasterizing the pixel grid as follows:
$\hat{S}_{i}(x,y)=\begin{cases}\mathbf{C}_{i}(x,y),&\text{if }M_{\mathrm{f}}(x,y)=1,\\ \mathbf{H}_{i}(x,y),&\text{if }M_{\mathrm{h}}(x,y)=1,\\ \mathbf{b},&\text{otherwise},\end{cases}$ (12)
where $\mathbf{b}\in\mathbb{R}^{3}$ is a predefined background color (e.g., black). This formulation ensures that observed pixels retain their original appearance, interior holes are smoothly inpainted from spatially proximate valid samples, and all remaining regions are filled with the background.
We integrate the point map with its RGB image to obtain a pose-aligned reprojected image, providing coarse 3D guidance for stitching. However, sparse inputs yield incomplete point clouds with holes and occlusions, leading to missing regions and misalignments in the reprojected image. To address this, we employ an image diffusion model to produce a visually coherent stitched result.
III-D Diffusion-based Stitching Refinement
@@
Figure 5: Qualitative comparison of different image restoration diffusion models. We evaluated the performance of several state-of-the-art image restoration diffusion models in refining reprojected images, including DiffIR [diffir_xia2023diffir], DDRM [ddrm_kawar2022denoising], and RDDM [liu2024residual].
III-E Implementation